A rabbit is 30 feet from its burrow. It can run 25 feet per second. A coyote that can run 50 feet per second spots the rabbit and starts running towards it. How can i write an expression and an equation to solve for the time it will take the rabbit to reach its burrow and the time for the coyote to get to the burrow?
Answer by stanbon(75887)

A rabbit is 30 feet from its burrow. It can run 25 feet per second. A coyote that can run 50 feet per second spots the rabbit and starts running towards it. How can i write an expression and an equation to solve for the time it will take the rabbit to reach its burrow and the time for the coyote to get to the burrow?
------------------
Rabbit DATA:
rate = 25 ft/sec ; distance = 30 ft; time = d/r = 30/25 = 6/5 seconds
----------------
Coyote DATA:
rate = 50 ft/sec ; distance = x ft ; time = d/r = x/50 seconds
------------------------
With this DATA you could determine the distance at which the coyote
could catch the rabbit before it reaches its burrow.
That equation would be x/50< 6/5
x < 60 ft
=============
Cheers,
Stan H.